OpenStudy (anonymous):

Which of the following is equivalent to 36ab^2 - 28ab when it is completely factored? A. 2ab(18b - 14) B. 4ab(9b - 7) C. 12ab(3b - 2) D. 4(9ab2 - 7ab)

OpenStudy (campbell_st):

well what are the common factors of 36ab^2 and 28ab or to make it easier to see \[4 \times 9 \times a \times b \times b... and.... 4 \times - 7 \times a \times b\] the common factors are outside the brackets and everything left behind is inside

OpenStudy (anonymous):

So is it A B C or D

OpenStudy (anonymous):

It's B. Try to learn this next time man.

OpenStudy (campbell_st):

well as @thefoe1990 said... try and learn it... it just makes life easier

OpenStudy (anonymous):

Crap I didn't even see you trying to help him. Sorry about that.